It takes one hour longer to fly to St. Paul at 200 mi/h than it does to return at 250 mi/h. How far away is St. Paul? I think it involves the use of the distance formula d=rt but I'm not sure how to use it because this problem only gives me "one hour longer" as a time.
Answer by stanbon(75887)

It takes one hour longer to fly to St. Paul at 200 mi/h than it does to return at 250 mi/h.
---------------------
Slow leg:
rate = 200 mph ; time = x hrs. : distance = rt = 200x miles
-----------------------
Fast leg:
rate = 250 mph ; time = x-1 hrs ; distance = rt = 250(x-1) miles
---------------------------
Equation:
distance = distance
200x = 250(x-1)
200x = 250x - 250
-50x = -250
x = 5 hrs.
---
Distance = 200x = 200*5 = 1000 miles
========================================
Cheers,
Stan H.






How far away is St. Paul? I think it involves the use of the distance formula d=rt but I'm not sure how to use it because this problem only gives me "one hour longer" as a time.